OpenStudy (anonymous):

evaluate for x=3, y=2, and z=1 (yz)^2

OpenStudy (anonymous):

i think it is 4 not sure

OpenStudy (anonymous):

its 4

OpenStudy (anonymous):

but how do i get it

OpenStudy (anonymous):

multiply 1 times 2 and square it and you get four so 2 times 1 is 2 square it its 4

OpenStudy (itsbribro):

(2*1)^2= 2^2= 2*2= 4

OpenStudy (anonymous):

ohhhh so i would do [(2)(1)]^2 then i would so [2]^2 then i would get 4

OpenStudy (itsbribro):

right or you could simply do it the way i did too! but ya! :)

OpenStudy (anonymous):

thank you so much :)

OpenStudy (itsbribro):

your welcome

OpenStudy (itsbribro):

do you have anymore questions?

OpenStudy (anonymous):

not at this moment but thanks for your help :)